### v1

@@ I Introduction
In this work, we employ a recently developed lattice Boltzmann framework [59] to study a model system: the dynamics of prolate and oblate spheroids immersed in a viscous fluid and confined within a channel of square cross section. The particle is driven by an external magnetic [66] or gravitational field, which applies a force on the particle. We systematically investigate the roles of confinement and inertia on the translational dynamics of the spheroid. We find that for a given volume, the maximum translational velocity of an unconfined spheroid does not occur for a sphere but rather for a prolate or an oblate shape depending on its orientation relative to the applied force. In confinement, the optimal aspect ratio shifts toward oblate shapes, highlighting a promising direction for future experimental studies in drug delivery applications. Furthermore, analysis of trajectories reveals that spheroids display glancing–reversing motions depending on the initial conditions, aspect ratio, and degree of confinement. At finite Reynolds numbers, inertial nonlinearities alter this behaviour, and by systematically increasing the Reynolds number from the Stokes regime to $\mathcal{O}(1)$, we uncover the emergence of rich nonlinear dynamics in the motion of confined spheroids.
@@ II.2 Translation of the confined spheroid
Eqs. 8 - 18 can be used to determine approximately the coupled translational-rotational dynamics of the spheroid. The complete dynamics can be captured using numerical simulations as described in the next section.
@@ III.3 Dynamics of the spheroid
An implicit numerical scheme is used to update the translational and angular velocities of the spheroid from the force and torque determined in the previous section. In this scheme [45], the total force and torque on the particle are written as,
@@ III.4 Summary of the simulation method
The simulation parameters used in this study are summarised below. For the unconfined cases, a cubic computational domain of size $256^{3}$ with periodic boundary conditions applied on all boundaries is employed. The major and minor axes of the spheroid vary between 6606-60 and 262-6, respectively. To investigate the role of geometric confinement, simulations are carried out in a domain of size $256\times N_{y}\times N_{z}$, where $N_{y}=N_{z}$ is varied from 51545-154 to vary the strength of confinement. When a spheroid is in close proximity to a wall, a lubrication force—based on the sphere–wall interaction and evaluated at the surface separation corresponding to the point of closest approach—is added to prevent wall contact. Periodic boundary conditions are applied on the open sides of the domain. To examine the influence of inertia, the kinematic viscosity of the fluid is varied 0.0070.50.007-0.5. In all cases, simulations are initialised with a quiescent fluid. Each run is continued until a steady state is reached, and the corresponding results are presented below.
@@ IV Results and Discussion
Figure 2: Translational velocity of an unconfined spheroid as a function of aspect ratio for a fixed particle volume. Aspect ratios $b/a<1$ correspond to prolate spheroids, while $b/a>1$ represent oblate shapes. The continuous curves show the analytical predictions (Sec. II.1), the symbols denote the lattice Boltzmann simulation results, and the dashed lines indicate the asymptotic expressions for rod-like and disc-like limits. Data plotted in blue (red) correspond to end-on (broadside-on) configurations, respectively. The aspect ratio at which the translational velocity is maximum in each case is marked by a vertical line.
@@ IV.1 Optimum aspect ratio of an unconfined spheroid
Under the action of the external force $\mathbf{F}_{e}$, the spheroid accelerates, but the drag force generated by the fluid opposes this motion. Balancing the two opposing forces, the particle undergoes a steady translation with a constant velocity. The variation in the translational velocity of unconfined prolate and oblate spheroids in an otherwise quiescent fluid as a function of aspect ratio $b/a$ is shown in Fig. 2.
@@ IV.2 Confined spheroid: centreline of the channel
In this section, we investigate the influence of confinement on the translational velocity of a spheroid by placing it inside a channel with a square cross‐section. As illustrated in Fig. 1(a), the external force is applied along the open direction of the channel (the $x$-axis). For simplicity, the spheroid is positioned on the channel centreline in an end‐on configuration, i.e., its axis $\hat{\bm{e}}$ is aligned with the channel axis.
@@ IV.3 Confined spheroid: eccentric positions in the channel
Figure 5: Schematic representation of (a) glancing and (b) reversing spheroid. The dotted line represents the channel centreline. The instantaneous orientation, indicated by the arrow $\hat{\bm{e}}$, is provided as a visual guide to illustrate the angular rotation of the spheroid. Glancing (reversing) trajectories span the full (half) channel width, with the spheroid approaching each wall almost parallel (perpendicular) to it. (c)-(d) Velocity vectors in the $h-\theta$ phase space for a spheroid of aspect ratio $b/a\approx 0.36$ and $b/a\approx 0.7$. The background is coloured with the angular velocity. The fixed points are highlighted with different markers (see text). In each figure, the inaccessible areas due to the finite size of the spheroid is coloured in grey.
@@ IV.5 Effect of fluid inertia and resulting bifurcations
In the results discussed so far, we have neglected the effects of fluid inertia and assumed the Stokes regime. We now relax this assumption and examine the influence of fluid inertia on the dynamics of a channel-confined spheroid. We fix the aspect ratio of the prolate spheroid at $b/a\approx 0.51$ and vary the Reynolds number $\mathcal{R}e$}. Note that the quasi-static simulations cannot be used to study the effect of inertia; we obtain the complete trajectory of the spheroid using LBM simulations.
@@ Acknowledgements
  • [3] M. N. Ardekani, P. Costa, W. P. Breugem, and L. Brandt (2016) Numerical study of the sedimentation of spheroidal particles. Int. J. Multiphase Flow 87, pp. 16–34. Cited by: §I.
@@
  • [15] V. Dabade, N. K. Marath, and G. Subramanian (2015) Effects of inertia and viscoelasticity on sedimenting anisotropic particles. J. Fluid Mech. 778, pp. 133–188. Cited by: §I, §IV.5.
@@
  • [20] J. Feng, H. H. Hu, and D. D. Joseph (1994) Direct simulation of initial value problems for the motion of solid bodies in a newtonian fluid part 1. sedimentation. J. Fluid Mech. 261, pp. 95–134. Cited by: §I, §IV.5.
@@
  • [29] P. Huang, H. H. Hu, and D. D. Joseph (1998) Direct simulation of the sedimentation of elliptic particles in oldroyd-b fluids. J. Fluid Mech. 362, pp. 297–326. Cited by: §I, §I, §IV.5.
@@
  • [34] R. Khayat and R. Cox (1989) Inertia effects on the motion of long slender bodies. J. Fluid Mech. 209, pp. 435–462. Cited by: §I, §IV.5.
@@
  • [42] W. H. Mitchell and S. E. Spagnolie (2015) Sedimentation of spheroidal bodies near walls in viscous fluids: glancing, reversing, tumbling and sliding. J. Fluid Mech. 772, pp. 600–629. Cited by: §I, §I, §I, §II.2, §IV.4.
@@
  • [50] T. Pan, R. Glowinski, and G. P. Galdi (2002) Direct simulation of the motion of a settling ellipsoid in newtonian fluid. J. Comput. Appl. Math. 149 (1), pp. 71–82. Cited by: §I, §I, §IV.5.

### v2

@@ I Introduction
In this work, we employ a recently developed lattice Boltzmann framework [61] to study a model system: the dynamics of prolate and oblate spheroids immersed in a viscous fluid and confined within a channel of square cross section. The particle is driven by an external magnetic [68] or gravitational field, which applies a force on the particle. We systematically investigate the roles of confinement and inertia on the translational dynamics of the spheroid. We consider only driven particles (by an external force) in this work which are experimentally easy to realise, say, compared to active particles for which the leading force description is a force dipole. We find that for a given volume, the maximum translational velocity of an unconfined spheroid does not occur for a sphere but rather for a prolate or an oblate shape depending on its orientation relative to the applied force. In confinement, the optimal aspect ratio shifts toward oblate shapes, highlighting a promising direction for future experimental studies in drug delivery applications. Furthermore, analysis of trajectories reveals that spheroids display glancing–reversing motions depending on the initial conditions, aspect ratio, and degree of confinement. At finite Reynolds numbers, inertial nonlinearities alter this behaviour, and by systematically increasing the Reynolds number from the Stokes regime to $\mathcal{O}(1)$, we uncover the emergence of rich nonlinear dynamics in the motion of confined spheroids.
@@ II Theory
The force $\bm{F}_{e}$ may be any body force applied externally and acts on the particle to drive the particle to translational motion. For example, the external force field can be considered to arise from a magnetic field when the spheroid is a microrobot, or gravitational field when gravitational sedimentation is considered.
We consider the translation of unconfined and confined spheroids. Here, ‘unconfined’ refers to the translation of the spheroids in the bulk fluid where wall effects are not present. ‘Confined’ spheroids refer to particles that are enclosed within the four walls of the square channel, as shown in Fig. 1(a), and thus the hydrodynamic interactions with the walls play a dominant role in their dynamics. The strength of the confinement is described by defining a confinement ratio, CR = $2b/L$ where $L$ is the height (equal to the depth) of the channel.
@@ II.2 Translation of the confined spheroid
Analysing Eqs. 10-12 we can see that the far-field prediction is $\mathcal{O}((a/h)^{3})$ for translational and $\mathcal{O}((a/h)^{4})$ for angular velocity. Eqs. 8 - 18 can be used to determine approximately the coupled translational-rotational dynamics of the spheroid. The complete dynamics can be captured using numerical simulations as described in the next section.
@@ III.3 Dynamics of the spheroid
An implicit numerical scheme is used to update the translational and angular velocities of the spheroid from the force and torque determined in the previous section, accounting for particle inertia. In this scheme [46], the total force and torque on the particle are written as,
@@ III.4 Summary of the simulation method
The simulation parameters used in this study are summarised below. For the unconfined cases, a cubic computational domain of size $256^{3}$ with periodic boundary conditions applied on all boundaries is employed. The major and minor axes of the spheroid vary between 6606-60 and 262-6, respectively. We have used particles with density ratio $\rho_{p}/\rho_{f}=1$ in the simulations with an external body force specified using $\bm{F}_{e}$. To investigate the role of geometric confinement, simulations are carried out in a domain of size $256\times N_{y}\times N_{z}$, where $N_{y}=N_{z}$ is varied from 51545-154 to vary the strength of confinement. When a spheroid is in close proximity to a wall, the hydrodynamic lubrication between the two can be under-represented on the discrete lattice. In this case a correction to the lubrication is made based on the surface separation at the point of closest approach. This prevents the spheroid touching the wall. Our results are independent of the specific implementation of this contact force. Periodic boundary conditions are applied on the open sides of the domain. To examine the influence of inertia, the kinematic viscosity of the fluid is varied 0.0070.50.007-0.5. In all cases, simulations are initialised with a quiescent fluid. Each run is continued until a steady state is reached, and the corresponding results are presented below.
@@ IV Results and Discussion
Figure 2: Steady state translational velocity of an unconfined spheroid as a function of aspect ratio for a fixed particle volume. Aspect ratios $b/a<1$ correspond to prolate spheroids, while $b/a>1$ represent oblate shapes. The continuous curves show the analytical predictions (Sec. II.1), the symbols denote the lattice Boltzmann simulation results, and the dashed lines indicate the asymptotic expressions for rod-like and disc-like limits. In the simulations, the spheroids were initially placed at the centre of the domain, with their symmetry axis either parallel ($\parallel$) or perpendicular ($\perp$) to the applied force ($\bm{F}_{e}=0.05$ lattice units). The spheroids retained this orientation through out the simulations. $\mathcal{R}e$}, calculated based on the steady state translational velocity of the spheroid, <<1<<1. For the sphere ($b/a=1$), ${\mathcal{R}e}\approx 0.023$. Data plotted in blue (red) correspond to end-on (broadside-on) configurations, respectively. The aspect ratio at which the translational velocity is maximum in each case is marked by a vertical line.
@@ IV.1 Optimum aspect ratio of an unconfined spheroid
Under the action of the external force $\bm{F}_{e}$, the spheroid accelerates, but the drag force generated by the fluid opposes this motion. Balancing the two opposing forces, the particle undergoes a steady translation with a constant velocity. The variation in the translational velocity of unconfined prolate and oblate spheroids in an otherwise quiescent fluid as a function of aspect ratio $b/a$ is shown in Fig. 2.
@@ IV.2 Confined spheroid: centreline of the channel
In this section, we investigate the influence of confinement on the translational velocity of a spheroid by placing it inside a channel with a square cross‐section. As illustrated in Fig. 1(a), the external force is applied along the open direction of the channel (the $x$-axis). For simplicity, the spheroid is positioned on the channel centreline in an end‐on configuration, i.e., its axis $\hat{\bm{e}}$ is aligned with the channel axis. As in the previous section, we apply a constant body force density to the spheroid by keeping the volume of the spheroid and the applied body force $\bm{F}_{e}$ constant while changing the aspect ratio and the confinement ratio.
@@ IV.3 Confined spheroid: eccentric positions in the channel
Figure 5: Schematic representation of (a) glancing and (b) reversing spheroid. The dotted line represents the channel centreline. The instantaneous orientation, indicated by the arrow $\hat{\bm{e}}$, is provided as a visual guide to illustrate the angular rotation of the spheroid. Glancing (reversing) trajectories span the full (half) channel width, with the spheroid approaching each wall almost parallel (perpendicular) to it. (c)-(d) Velocity vectors in the $h-\theta$ phase space for a spheroid of aspect ratio $b/a\approx 0.36$ and $b/a\approx 0.7$. $h$ is normalised by the channel width $L$ and $\theta$ is given in degrees (). The background is coloured with the angular velocity; $\varOmega_{z}>0$ corresponds to anti-clockwise rotation. Black vector arrows indicate the direction of the velocity of the spheroid in the phase space. The fixed points are highlighted with different markers (see text). In each figure, the inaccessible areas due to the finite size of the spheroid is coloured in grey.
@@ IV.5 Effect of fluid inertia and resulting bifurcations
In the results discussed so far, we have neglected the effects of fluid inertia and assumed the Stokes regime. We now relax this assumption and examine the influence of fluid inertia on the dynamics of a channel-confined spheroid. We fix the aspect ratio of the prolate spheroid at $b/a\approx 0.51$ and vary the Reynolds number $\mathcal{R}e$}. The reported $\mathcal{R}e$} is based on the steady state velocity of the translating spheroid. Note that the quasi-static simulations cannot be used to study the effect of inertia; we obtain the complete trajectory of the spheroid using LBM simulations.
@@
The presence of spiralling trajectories in the phase space indicates that fluid inertia also breaks the symmetry in the translational and angular kinematics of the spheroid with respect to the channel centreline. The extent of asymmetry is proportional to the Reynolds number $\mathcal{R}e$}. We find that both components of translational velocity and the angular velocity, $U_{x}$, $U_{y}$ and $\Omega_{z}$ are visibly not symmetric with respect to the channel centreline when quasi-static simulations are performed at ${\mathcal{R}e}\approx\mathcal{O}(0.1)$. With regard to the final configuration of the particle, previous studies have shown that fluid-inertia-induced torque on non-spherical particles cannot be neglected even at Reynolds numbers much smaller than unity [15, 58]. The preferred orientation of a translating spheroid due to finite inertia is at $\theta=90^{\circ}$, i.e., the broadside-one configuration [35, 29, 51, 20, 15, 32, 3]. Our results show that, even in confinement this result holds; but the location of the equilibrium positions can be either at the channel centre line or near the walls depending upon $\mathcal{R}e$} as shown in Fig. 7. Hence, the approximate results obtained by Mitchell and Spagnolie [43] which account for the wall induced hydrodynamic interaction can be substantially altered due to the presence of a small but finite Reynolds number.
@@ V Conclusions
Conflicts of interest
There are no conflicts of interest to declare.
@@ Acknowledgements
  • [3] M. N. Ardekani, P. Costa, W. P. Breugem, and L. Brandt (2016) Numerical study of the sedimentation of spheroidal particles. Int. J. Multiphase Flow 87, pp. 16–34. Cited by: §I, §IV.5.
@@
  • [15] V. Dabade, N. K. Marath, and G. Subramanian (2015) Effects of inertia and viscoelasticity on sedimenting anisotropic particles. J. Fluid Mech. 778, pp. 133–188. Cited by: §I, §IV.5, §IV.5.
@@
  • [20] J. Feng, H. H. Hu, and D. D. Joseph (1994) Direct simulation of initial value problems for the motion of solid bodies in a newtonian fluid part 1. sedimentation. J. Fluid Mech. 261, pp. 95–134. Cited by: §I, §IV.5, §IV.5.
@@
  • [29] P. Huang, H. H. Hu, and D. D. Joseph (1998) Direct simulation of the sedimentation of elliptic particles in oldroyd-b fluids. J. Fluid Mech. 362, pp. 297–326. Cited by: §I, §I, §IV.5, §IV.5.
@@
  • [32] X. Jiang, C. Xu, and L. Zhao (2024) Prolate spheroids settling in a quiescent fluid: clustering, microstructures and collisions. J. Fluid Mech. 1000, pp. A49. Cited by: §IV.5.
@@
  • [35] R. Khayat and R. Cox (1989) Inertia effects on the motion of long slender bodies. J. Fluid Mech. 209, pp. 435–462. Cited by: §I, §IV.5, §IV.5.
@@
  • [43] W. H. Mitchell and S. E. Spagnolie (2015) Sedimentation of spheroidal bodies near walls in viscous fluids: glancing, reversing, tumbling and sliding. J. Fluid Mech. 772, pp. 600–629. Cited by: §I, §I, §I, §II.2, §IV.4, §IV.5.
@@
  • [51] T. Pan, R. Glowinski, and G. P. Galdi (2002) Direct simulation of the motion of a settling ellipsoid in newtonian fluid. J. Comput. Appl. Math. 149 (1), pp. 71–82. Cited by: §I, §I, §IV.5, §IV.5.
@@
  • [58] M. Z. Sheikh, K. Gustavsson, D. Lopez, E. Lévêque, B. Mehlig, A. Pumir, and A. Naso (2020) Importance of fluid inertia for the orientation of spheroids settling in turbulent flow. Journal of Fluid Mechanics 886, pp. A9. Cited by: §IV.5.